# PAC Learnability in the Presence of Performativity

### v1

@@ 5 Learnability under Other Performative Maps
We assume the following conditions hold for all $h\in\mathcal{H}$: (i) the measure $\tilde{\mathbb{P}}_{h}$ is absolutely continuous with respect to $\mathbb{P}$; and (ii) the RN derivative is known to the learner and uniformly bounded—specifically, there exists a constant $M\geq 0$ such that for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, $\frac{d\tilde{\mathbb{P}}_{h}}{d\mathbb{P}}(x,y)\leq M$.
The first assumption is a standard condition ensuring the existence of the RN derivative. The second assumption is a counterpart to the learner knowing the distribution shift map in Section˜3.2.
@@
$\text{PR}(h_{S}^{\ast})\leq\min_{h\in\mathcal{H}}\text{PR}(h)$ $+4M\mathfrak{R}_{n}(\mathcal{H})$
$+4M(1+\sqrt{2})\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}.$
@@ Appendix C Proofs
$\text{PR}(h_{S}^{\ast})\leq\min_{h\in\mathcal{H}}\text{PR}(h)+4M\mathfrak{R}_{n}(\mathcal{H})+4M(1+\sqrt{2})\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}.$
@@
  • Step 3. Now we need to relate the Rademacher averages between the classes $\Omega_{\mathcal{H}}$ and $\mathcal{H}$. We have
    $\mathfrak{R}(\Omega_{\mathcal{H}})$ $=\mathbb{E}\bigg[\sup_{\omega_{h}\in\Omega_{\mathcal{H}}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\omega_{h}(X_{i},Y_{i})\bigg]$
    $=\mathbb{E}\bigg[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\frac{d\tilde{\mathbb{P}}_{h}}{d\mathbb{P}}(X_{i},Y_{i})\mathbf{1}_{h(X_{i})\neq Y_{i}}\bigg]$
    $\leq\mathbb{E}\bigg[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}M\mathbf{1}_{h(X_{i})\neq Y_{i}}\bigg]$
    $=M\mathfrak{R}(\mathcal{H})$
    where we use that the Radon-Nikodym derivative is bounded by $M$. Combining this with the inequality in Step 2, we get that with with probability at least $1-\delta/4$
    $\sup_{h\in\mathcal{H}}\{\text{PR}(h)-\text{PR}_{n}(h)\}\leq 2M\mathfrak{R}(\mathcal{H})+\sqrt{\frac{4M^{2}\ln{\frac{4}{\delta}}}{2n}},$
    which establishes the first part of the theorem.
@@
    $\sup_{h\in\mathcal{H}}\{\text{PR}(h)-\text{PR}_{n}(h)\}\leq 2M\mathfrak{R}_{n}(\mathcal{H})+2M(1+\sqrt{2})\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}},$
@@
The rest of the theorem is essentially the same as in Mohri et al. (2018) and we give it here for completeness only. We have that for every $h\in\mathcal{H}$, with probability at laest $1-\delta$
$\text{PR}(h_{S}^{\ast})$ $\leq\text{PR}_{n}(h_{S}^{\ast})+2M\mathfrak{R}_{n}(\mathcal{H})+2M(1+\sqrt{2})\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}$
$\leq\text{PR}_{n}(h)+2M\mathfrak{R}_{n}(\mathcal{H})+2M(1+\sqrt{2})\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}$
$\leq\text{PR}(h)+4M\mathfrak{R}_{n}(\mathcal{H})+4M(1+\sqrt{2})\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}.$

### v2

@@ 5 Learnability under Other Performative Maps
We assume the following conditions hold for all $h\in\mathcal{H}$: (i) the measure $\tilde{\mathbb{P}}_{h}$ is absolutely continuous with respect to $\mathbb{P}$; (ii) the RN derivative is uniformly bounded—specifically, there exists a constant $M\geq 0$ such that for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, $\frac{d\tilde{\mathbb{P}}_{h}}{d\mathbb{P}}(x,y)\leq M$; and (iii) the function $\frac{d\tilde{\mathbb{P}}_{h}}{d\mathbb{P}}$ is uniquely determined via the values of $h$, i.e. that $\frac{d\tilde{\mathbb{P}}_{h}}{d\mathbb{P}}(x,y)=\psi(x,y,h(x))$ for some function $\psi$. [^1] The third assumption was added in the second version of this manuscript, to fix an incorrect bounding argument in Step 3 of the initial proof of Theorem 5.1. We apologise for this unintentional mistake and thank an anonymous reviewer for detecting the issue in the previous version.
The first assumption is a standard condition ensuring the existence of the RN derivative. The second assumption bounds the strength of the performative effect. The third one states that the change in distribution only depends on the classifier via its values and not, say, on its parameters. Note that, as in Section 4, the learners we study below depend on the distribution map (via the RN derivative), as means towards proving the existence of a PAC learner in the sense of Definition 3.1.
@@
$\text{PR}(h_{S}^{\ast})\leq\min_{h\in\mathcal{H}}\text{PR}(h)$ $+2M\mathfrak{R}_{n}(\mathcal{H})$
$+(4+2\sqrt{2})M\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}.$
@@ Appendix C Proofs
$\text{PR}(h_{S}^{\ast})\leq\min_{h\in\mathcal{H}}\text{PR}(h)+2M\mathfrak{R}_{n}(\mathcal{H})+(4+2\sqrt{2})M\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}.$
@@
  • Step 3. Now we need to relate the Rademacher complexity of the class:
    $\Omega_{\mathcal{H}}=\left\{(x,y)\mapsto\frac{d\tilde{\mathbb{P}}_{h}}{d\mathbb{P}}(x,y)\cdot 1_{h(x)\neq y}\mid h\in\mathcal{H}\right\}.$
    to the Rademacher complexity of $\mathcal{H}$. We do so by first considering the empirical Rademacher complexity with respect to an i.i.d. sample $S=\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ from $\mathbb{P}$. For each index $i$, we define the function $\phi_{i}:\{-1,1\}\to\mathbb{R}$ as:
    $\phi_{i}(u)$ $:=\frac{d\tilde{\mathbb{P}}_{u}}{d\mathbb{P}}(x_{i},y_{i})\cdot 1_{u\neq y_{i}}$
    $=\psi(x_{i},y_{i},u)\cdot 1_{u\neq y_{i}}.$
    The empirical Rademacher complexity is given by:
    $\hat{\mathfrak{R}}_{S}(\Omega_{\mathcal{H}})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi_{i}(h(x_{i}))\right].$
    We seek to bound $\hat{\mathfrak{R}}_{S}(\Omega_{\mathcal{H}})$ by using Talagrand’s Lemma. To this end, we must determine the Lipschitz constant of $\phi_{i}$ with respect to its argument $u\in\{-1,1\}$. The Lipschitz condition requires $|\phi_{i}(1)-\phi_{i}(-1)|\leq L|1-(-1)|=2L$.
    We analyze the difference $|\phi_{i}(1)-\phi_{i}(-1)|$:
    • Case 1 ($y_{i}=1$): Here $1_{1\neq 1}=0$ and $1_{-1\neq 1}=1$.
      $|\phi_{i}(1)-\phi_{i}(-1)|=|0-\psi(x_{i},1,-1)|=\psi(x_{i},1,-1)\leq M.$
    • Case 2 ($y_{i}=-1$): Here $1_{1\neq-1}=1$ and $1_{-1\neq-1}=0$.
      $|\phi_{i}(1)-\phi_{i}(-1)|=|\psi(x_{i},-1,1)-0|=\psi(x_{i},-1,1)\leq M.$
    In both cases, the difference is bounded by $M$. Therefore, $2L=M\implies L=M/2$.
    Therefore, we can apply Talagrand’s Contraction Lemma. Specifically, Lemma 5.7 in Mohri et al. (2018) states that if $\phi_{i}$ are $L$-Lipschitz functions, then
    $\mathbb{E}_{\sigma}\left(\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi_{i}(h(x_{i}))\right)\leq L\mathbb{E}_{\sigma}\left(\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(x_{i})\right).$
    Applying this with $L=M/2$ gives:
    $\hat{\mathfrak{R}}_{S}(\Omega_{\mathcal{H}})\leq\frac{M}{2}\hat{\mathfrak{R}}_{S}(\mathcal{H}).$
    Finally, we take expectation with respect to the sample, to obtain a link between the distributional Rademacher complexities. Combining this with the inequality in Step 2, we get that with with probability at least $1-\delta/4$
    $\sup_{h\in\mathcal{H}}\{\text{PR}(h)-\text{PR}_{n}(h)\}\leq M\mathfrak{R}(\mathcal{H})+\sqrt{\frac{4M^{2}\ln{\frac{4}{\delta}}}{2n}}.$
@@
    $\sup_{h\in\mathcal{H}}\{\text{PR}(h)-\text{PR}_{n}(h)\}\leq M\mathfrak{R}_{n}(\mathcal{H})+(2+\sqrt{2})M\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}},$
@@
The rest of the theorem is essentially the same as in Mohri et al. (2018) and we give it here for completeness only. We have that for every $h\in\mathcal{H}$, with probability at least $1-\delta$
$\text{PR}(h_{S}^{\ast})$ $\leq\text{PR}_{n}(h_{S}^{\ast})+M\mathfrak{R}_{n}(\mathcal{H})+(2+\sqrt{2})M\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}$
$\leq\text{PR}_{n}(h)+M\mathfrak{R}_{n}(\mathcal{H})+(2+\sqrt{2})M\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}$
$\leq\text{PR}(h)+2M\mathfrak{R}_{n}(\mathcal{H})+(4+2\sqrt{2})M\sqrt{\frac{\ln{\frac{4}{\delta}}}{2n}}.$